OpenStudy (anonymous):

The express sqrt(180) minus sqrt(20) equals ? times sqrt(5)

OpenStudy (anonymous):

\[\sqrt{9*4*5}-\sqrt{4*5}=6\sqrt{5}-2\sqrt{5}=4\sqrt{5}\]